# Multiscale thermodynamic nonequilibrium effects in Kelvin-Helmholtz instability and their relative importance

### v2

@@ I Introduction
In recent years, DBM has been extensively applied to a wide range of complex fluid systems, including: liquid-vapor phase-separating systems Gan et al. (2015, 2022), hydrodynamic instabilitiesLai et al. (2016); Chen et al. (2025, 2024); Xu, Lin, and Lai (2025), reactive flows (e.g., detonation and combustion) Zhang et al. (2016); Lin et al. (2016a); Ji, Lin, and Luo (2022), Shock reflection and shock wave/boundary layer interaction Bao et al. (2022); Qiu et al. (2020, 2021), micro-nanoscale flowsZhang et al. (2019, 2023), plasma kineticsLiu et al. (2023); Song et al. (2024), etc. As an analytical tool for complex physical fields, DBM can be used to characterize the evolution stages of complex flows and analyze the physical mechanisms of flow systems. For instance, the temporal evolution of non-equilibrium intensity $D$ has been employed to distinguish different stages of phase separation, including the spinodal decomposition stage, the transition to domain growth, and the domain growth stage Gan et al. (2015). Similarly, the evolution of the non-equilibrium component $\Delta_{3,1y}^{*}$ has been used to characterize various stages of RTI development and serve as a physical marker for interface tracking in complex flow simulations Lai et al. (2016). The evolution of viscous stress has been utilized to classify different stages of bubble coalescence Sun et al. (2024), etc.
@@
In recent years, significant progress has also been made in studying KHI and its nonequilibrium effects using DBM Gan et al. (2011); Lin et al. (2016b); Gan et al. (2019); Lin et al. (2019, 2021); Li and Lin (2024).
@@
Early foundations: In 2011, Gan et al. employed the lattice Boltzmann kinetic method, a precursor to DBM, to investigate the competing effects of the velocity transition layer ($L_{u}$) and the density transition layer ($L_{\rho}$) on KHI growthGan et al. (2011). They introduced the dimensionless parameter $R=L_{\rho}/L_{u}$ to quantify these interactions, which was later validated through theoretical analysisWang, Ye, and Li (2010). Building on this, Lin et al. (2016) developed a compressible two-component DBM capable of capturing detailed interface deformations during KHI evolution, significantly improving multiscale flow modeling Lin et al. (2016b).
@@
Investigations into nonequilibrium effects: In 2019, Gan et al. applied DBM to examine the influence of viscosity, heat conduction, and the Prandtl number on KHI evolutionGan et al. (2019). They introduced a novel approach combining nonequilibrium behavior tracking with morphological analysis. This method enabled precise characterization of characteristic structures, as well as the width and growth rate of the KH mixing layer. Key findings included: (i) Strong spatiotemporal correlations between the mixing layer width, nonequilibrium intensity, and interface boundary length. (ii) Viscous stress suppresses KHI growth, prolongs the linear stage, reduces maximum disturbance energy, enhances both global and local thermodynamic nonequilibrium (TNE), and broadens the nonequilibrium region. (iii) Heat conduction inhibits small-wavelength growth, favoring the formation of large-scale structures. Under high heat conduction, an initial suppression phase is followed by enhancement, with suppression dominating early stages and enhancement prevailing later. That same year, Lin et al. investigated the role of relaxation time in nonequilibrium behaviors during KHI evolution. Their results indicated that longer relaxation times increased global nonequilibrium intensity, mixing entropy, and mixing free energy. For smaller Atwood numbers, nonequilibrium peaked earlier, with rapid changes in mixing entropy and free energyLin et al. (2019).
@@
Despite significant progress in applying DBM to KHI studies, these models face severe challenges under extreme conditions. Such conditions include increased shear rates, narrowing of velocity, material, and mechanical interfaces, the emergence of additional nonequilibrium driving forces, the presence of multiple fluid instabilities, and enhanced nonlinear coupling effects. Under these circumstances, KHI exhibits persistent and pronounced TNEs and hydrodynamic nonequilibrium effects (HNEs), posing substantial challenges to model validity. Extreme flow environments frequently give rise to such conditions. For instance, in ICF, laser or heavy-ion beams induce steep density and velocity gradients at the shell interface, leading to rapid KHI development. In this regime, strong nonequilibrium effects dominate interface evolution and mixing behaviorHe and Zhang (2007); Soler and Ballester (2022). In astrophysics, KHI rapidly develops during supernova explosions and stellar wind interactions with interstellar media. These processes produce complex vortex and turbulence structures due to sharp fluid interfaces and strong velocity shearBurrows (2000); Drake (2010); Hasegawa et al. (2004). Similarly, in the wake of high-speed aircraft, supersonic jets experience rapid KHI growth driven by strong shear layers, where nonequilibrium effects critically impact wake structures and flow evolutionMartens, Kinzie, and McLaughlin (1994); Shi et al. (2023).
@@ II Discrete Boltzmann modeling and complex physical field analysis method
$\partial_{t}(\rho e)+\bm{\nabla}\cdot\left[(e+P)\mathbf{u}+\bm{\Delta}_{2}^{*}\cdot\mathbf{u}+\bm{\Delta}_{3,1}^{*}\right]=0.$ (8)
@@ III Key considerations for constructing discrete velocity sets
As a theoretical framework for model construction and complex physical field analysis, the DBM imposes only the essential physical constraints required to describe the considered TNEs. It does not specify temporal, spatial, or phase-space discretization schemes. However, numerically, the choice of DVS plays a crucial role in determining the ability to accurately capture physical effects in DBM simulation. In 2018, Gan et al. proposed a two-dimensional discrete velocity set with 26 discrete velocities (D2V26) for phase-space discretization Gan et al. (2018). Building on this foundation, four new velocity sets (DVS1-DVS4) are developed to incorporate second-order nonequilibrium effects while optimizing computational efficiency and stability (Figure 1). The key characteristics of these sets are as follows: (i) DVS1: Retains the first 25 discrete velocities from Ref. Gan et al., 2018. (ii) DVS2: A more symmetric 25-velocity set designed for improved performance. (iii) DVS3: Extends DVS2 by introducing additional velocity magnitudes, resulting in a more diverse and evenly distributed velocity set. (iv) DVS4: Further optimizes velocity distribution by increasing the diversity of magnitude combinations. The detailed configurations of the four DVSs are given as follows:
@@
where $L_{\rho}$ and $L_{u}$ denote the widths of the density and velocity transition layers, respectively. The parameters $\rho_{L}$ and $\rho_{R}$ represent the fluid densities far from the interface on the left and right sides. The computational domain is a square with a side length of $0.3$, discretized using a uniform $200\times 200$ grid. The initial physical conditions are set as $204\rho_{L}=\rho_{R}=4,4P_{L}=P_{R}=4,u_{y}=0,L_{\rho}=L_{u}=$ 20, with parameters $\tau=2\times 10^{-3},u_{0}=1.5$, $\Delta t=1\times 10^{-5},\Delta x=\Delta y=0.0015$.
@@ IV Multiscale TNEs and their relative importance in KHI
Boundary conditions for the KHI simulations are implemented as follows: periodic boundary conditions are applied in the $y$-direction, while outflow boundary conditions, commonly used in KHI studies, are enforced in the $x$-direction Keppens and Tóth (1999); Perucho et al. (2007); Lifeng et al. (2009). Boundary conditions play a critical role in fluid models, directly influencing nonequilibrium behaviors. Properly designed kinetic boundary conditions are essential for ensuring physical accuracy, numerical stability, and computational precision.
@@ IV.1 First-order TNEs dominance case
Figure 4: Density evolution patterns of KHI at four time instances for the first-order TNEs dominance case: (a) $t=0.24$, (b) $t=0.60$, (c) $t=1.68$, (d) $t=2.40$.
@@ IV.2 Activation of second-order TNEs
Figure 7: Density evolution patterns of KHI at four time instances with the activation of second-order TNEs: (a) $t=0.24$, (b) $t=0.72$, (c) $t=1.68$, (d) $t=2.40$.
@@ IV.2.1 Comparison of viscous stress and its components
Figure 8 compares the viscous stress obtained from high-order (right column) and low-order (left column) models at various stages of KHI evolution, with the analytical solution. Compared to the weak nonequilibrium case, the deviation of the D2V16 results from the analytical solution increases significantly. In contrast, the D2V25 results closely align with the second-order nonequilibrium analytical solution, demonstrating its capability to capture stronger nonequilibrium effects.
@@
$T_{\rho 1}=-2n_{2}^{-2}\tau^{2}R^{2}T^{2}\left[n_{1}n_{2}\frac{\partial^{2}}{\partial x^{2}}\rho-n_{2}\frac{\partial^{2}}{\partial y^{2}}\rho\right].$ (30)
@@ IV.2.4 Comparison of macroscopic quantities
Figure 15: Density evolution patterns of the KHI at four time instances under the dominance of second-order TNEs: (a) $t=0.24$, (b) $t=0.72$, (c) $t=2.70$, and (d) $t=3.30$.
@@ IV.3 Dominance of second-order TNEs
providing a comprehensive characterization, we analyze the relative nonequilibrium strength: $R_{\text{TNE}}={\bm{\Delta}_{2}^{*(2)}}/{\bm{\Delta}_{2}^{*(1)}}$. The analytical expressions for $\Delta_{2xx}^{*(1)}$ and $\Delta_{2xx}^{*(2)}$ are provided in Eqs. 48 and 32. Dividing each term in $\Delta_{2xx}^{*(2)}$ by $\Delta_{2xx}^{*(1)}$ yields:
@@ IV.4.1 Absolute dominance of second-order TNEs
Figure 18: Density evolution patterns of the KHI at four time instances under the absolute dominance of second-order TNEs: (a) $t=0.18$, (b) $t=0.36$, (c) $t=0.54$, (d) $t=1.08$.
@@ V Conclusions, discussions, and outlooks
When second-order nonequilibrium effects come to the forefront, the battle between low- and high-order models becomes evident. Our comparative analysis reveals critical shortcomings in low-order models: they fall to accurately predict KHI growth rates, capture macroscopic evolution, preserve kinetic moment relationships, and faithfully reproduce thermodynamic nonequilibrium characteristics. In contrast, high-order models emerge as the superior contenders, offering sharper physical accuracy, greater cross-scale adaptability, and enhanced numerical stability in the presence of strong nonequilibrium effects. Notably, although high-order models exhibit some degree of extensibility in capturing strong nonequilibrium effects, their applicability remains fundamentally constrained. In physical modeling, the key lies in dynamically adjusting the preserved kinetic moment relations during phase-space discretization and tailoring discretization strategies to specific research objectives-only then can we ensure a precise and robust kinetic description. To further unravel the multiscale nature of KHI, we construct a nonequilibrium phase diagram-a roadmap that visually charts the interplay between different nonequilibrium orders. This diagram not only deepens our understanding of the instability’s intricate dynamics but also serves as a valuable guide for selecting suitable kinetic models across a broad spectrum of nonequilibrium conditions in both scientific research and engineering applications.
@@ Appendix B Nonlinear constitutive relations
$\bm{\Delta}^{*(1)}_{2}=-\mu\left[\bm{\nabla}u+(\bm{\nabla}u)^{T}-\frac{2}{n+2}\bm{\nabla}\cdot\mathbf{u}\right],$ (48)

### v3

@@ I Introduction
In recent years, DBM has been extensively applied to a wide range of complex fluid systems, including: liquid-vapor phase-separating systems Gan et al. (2015, 2022), hydrodynamic instabilitiesLai et al. (2016); Chen et al. (2025a, 2024); Xu, Lin, and Lai (2025), reactive flows (e.g., detonation and combustion) Zhang et al. (2016); Ji, Lin, and Luo (2022), Shock reflection and shock wave/boundary layer interaction Bao et al. (2022); Qiu et al. (2020, 2021), micro-nanoscale flowsZhang et al. (2019, 2023), plasma kineticsLiu et al. (2023); Song et al. (2024), etc. As an analytical tool for complex physical fields, DBM can be used to characterize the evolution stages of complex flows and analyze the physical mechanisms of flow systems. For instance, the temporal evolution of non-equilibrium intensity $D$ has been employed to distinguish different stages of phase separation, including the spinodal decomposition stage, the transition to domain growth, and the domain growth stage Gan et al. (2015). Similarly, the evolution of the non-equilibrium component $\Delta_{3,1y}^{*}$ has been used to characterize various stages of RTI development and serve as a physical marker for interface tracking in complex flow simulations Lai et al. (2016). The evolution of viscous stress has been utilized to classify different stages of bubble coalescence Sun et al. (2024), etc.
@@
In recent years, significant progress has also been made in studying KHI and its nonequilibrium effects using DBM Gan et al. (2011); Lin et al. (2016); Gan et al. (2019); Lin et al. (2019, 2021); Li and Lin (2024).
@@
Early foundations: In 2011, Gan et al. employed the lattice Boltzmann kinetic method, a precursor to DBM, to investigate the competing effects of the velocity transition layer ($L_{u}$) and the density transition layer ($L_{\rho}$) on KHI growthGan et al. (2011). They introduced the dimensionless parameter $R=L_{\rho}/L_{u}$ to quantify these interactions, which was later validated through theoretical analysisWang, Ye, and Li (2010). Building on this, Lin et al. (2016) developed a compressible two-component DBM capable of capturing detailed interface deformations during KHI evolution, significantly improving multiscale flow modeling Lin et al. (2016).
@@
Investigations into nonequilibrium effects: In 2019, Gan et al. applied DBM to examine the influence of viscosity, heat conduction, and the Prandtl number on KHI evolutionGan et al. (2019). They introduced a novel approach combining nonequilibrium behavior tracking with morphological analysis. This method enabled precise characterization of characteristic structures, as well as the width and growth rate of the KH mixing layer. Key findings included: (i) Strong spatiotemporal correlations between the mixing layer width, nonequilibrium intensity, and interface boundary length. (ii) Viscous stress suppresses KHI growth, prolongs the linear stage, reduces maximum disturbance energy, enhances both global and local thermodynamic nonequilibrium, and broadens the nonequilibrium region. (iii) Heat conduction inhibits small-wavelength growth, favoring the formation of large-scale structures. Under high heat conduction, an initial suppression phase is followed by enhancement, with suppression dominating early stages and enhancement prevailing later. That same year, Lin et al. investigated the role of relaxation time in nonequilibrium behaviors during KHI evolution. Their results indicated that longer relaxation times increased global nonequilibrium intensity, mixing entropy, and mixing free energy. For smaller Atwood numbers, nonequilibrium peaked earlier, with rapid changes in mixing entropy and free energyLin et al. (2019).
@@
Despite significant progress in applying DBM to KHI studies, these models face severe challenges under extreme conditions. Such conditions include increased shear rates, narrowing of velocity, material, and mechanical interfaces, the emergence of additional nonequilibrium driving forces, the presence of multiple fluid instabilities, and enhanced nonlinear coupling effects. Under these circumstances, KHI exhibits persistent and pronounced TNEs and HNEs, posing substantial challenges to model validity. Extreme flow environments frequently give rise to such conditions. For instance, in ICF, laser or heavy-ion beams induce steep density and velocity gradients at the shell interface, leading to rapid KHI development. In this regime, strong nonequilibrium effects dominate interface evolution and mixing behaviorHe and Zhang (2007); Soler and Ballester (2022). In astrophysics, KHI rapidly develops during supernova explosions and stellar wind interactions with interstellar media. These processes produce complex vortex and turbulence structures due to sharp fluid interfaces and strong velocity shearBurrows (2000); Drake (2010); Hasegawa et al. (2004). Similarly, in the wake of high-speed aircraft, supersonic jets experience rapid KHI growth driven by strong shear layers, where nonequilibrium effects critically impact wake structures and flow evolutionMartens, Kinzie, and McLaughlin (1994); Shi et al. (2023).
@@ II Discrete Boltzmann modeling and complex physical field analysis method
$\partial_{t}(\rho e)+\bm{\nabla}\cdot\left[(\rho e+P)\mathbf{u}+\bm{\Delta}_{2}^{*}\cdot\mathbf{u}+\bm{\Delta}_{3,1}^{*}\right]=0.$ (8)
@@ III Key considerations for constructing discrete velocity sets
As a theoretical framework for model construction and complex physical field analysis, the DBM imposes only the essential physical constraints required to describe the considered TNEs. It does not specify temporal, spatial, or phase-space discretization schemes. However, numerically, the choice of DVS plays a crucial role in determining the ability to accurately capture physical effects in DBM simulation. In 2018, Gan et al. proposed a two-dimensional discrete velocity set with 26 discrete velocities (D2V26) for phase-space discretization Gan et al. (2018). Building on this foundation, four new velocity sets (DVS1-DVS4) are developed to incorporate second-order nonequilibrium effects while optimizing computational efficiency and stability (Figure 1). The key characteristics of these sets are as follows: (i) DVS1: Retains the first 25 discrete velocities from Ref. Gan et al., 2018. (ii) DVS2: A more symmetric 25-velocity set designed for improved performance Chen et al. (2025b). (iii) DVS3: Extends DVS2 by introducing additional velocity magnitudes, resulting in a more diverse and evenly distributed velocity set. (iv) DVS4: Further optimizes velocity distribution by increasing the diversity of magnitude combinations. The detailed configurations of the four DVSs are given as follows:
@@
where $L_{\rho}$ and $L_{u}$ denote the widths of the density and velocity transition layers, respectively. The parameters $\rho_{L}$ and $\rho_{R}$ represent the fluid densities far from the interface on the left and right sides. The computational domain is a square with a side length of $0.3$, discretized using a uniform $200\times 200$ grid. The initial physical conditions are set as $204\rho_{L}=\rho_{R}=4,4P_{L}=P_{R}=4,u_{y}=0,L_{\rho}=L_{u}=$ 20, with parameters $\tau=2\times 10^{-3},u_{0}=1.5$, $\Delta t=10^{-5}$, $\Delta x=\Delta y=0.0015$.
@@ IV Multiscale TNEs and their relative importance in KHI
Boundary conditions for the KHI simulations are implemented as follows: periodic boundary conditions are applied in the $y$-direction, while outflow boundary conditions, commonly used in KHI studies, are enforced in the $x$-direction Keppens and Tóth (1999); Perucho et al. (2007); Wang et al. (2009). Boundary conditions play a critical role in fluid models, directly influencing nonequilibrium behaviors. Properly designed kinetic boundary conditions are essential for ensuring physical accuracy, numerical stability, and computational precision.
@@ IV.1 First-order TNEs dominance case
Figure 4: Density evolution patterns of KHI at four moments for the first-order TNEs dominance case: (a) $t=0.24$, (b) $t=0.60$, (c) $t=1.68$, (d) $t=2.40$.
@@ IV.2 Activation of second-order TNEs
Figure 7: Density evolution patterns of KHI at four moments with the activation of second-order TNEs: (a) $t=0.24$, (b) $t=0.72$, (c) $t=1.68$, (d) $t=2.40$.
@@ IV.2.1 Comparison of viscous stress and its components
$T_{\rho 1}=-2n_{2}^{-2}\tau^{2}R^{2}T^{2}\left(n_{1}n_{2}\frac{\partial^{2}}{\partial x^{2}}\rho-n_{2}\frac{\partial^{2}}{\partial y^{2}}\rho\right)$ (30)
@@ IV.2.4 Comparison of macroscopic quantities
Figure 15: Density evolution patterns of the KHI at four moments under the dominance of second-order TNEs: (a) $t=0.24$, (b) $t=0.72$, (c) $t=2.70$, and (d) $t=3.30$.
@@ IV.3 Dominance of second-order TNEs
To provide a comprehensive characterization, we analyze the relative nonequilibrium strength: $R_{\text{TNE}}={\bm{\Delta}_{2}^{*(2)}}/{\bm{\Delta}_{2}^{*(1)}}$. The analytical expressions for $\Delta_{2xx}^{*(1)}$ and $\Delta_{2xx}^{*(2)}$ are provided in Eqs. 48 and 32. Dividing each term in $\Delta_{2xx}^{*(2)}$ by $\Delta_{2xx}^{*(1)}$ yields:
@@ IV.4.1 Absolute dominance of second-order TNEs
Figure 18: Density evolution patterns of the KHI at four moments under the absolute dominance of second-order TNEs: (a) $t=0.18$, (b) $t=0.36$, (c) $t=0.54$, (d) $t=1.08$.
@@ V Conclusions, discussions, and outlooks
When second-order nonequilibrium effects come to the forefront, the battle between low- and high-order models becomes evident. Our comparative analysis reveals critical shortcomings in low-order models: they fail to accurately predict KHI growth rates, capture macroscopic evolution, preserve kinetic moment relationships, and faithfully reproduce thermodynamic nonequilibrium characteristics. In contrast, high-order models emerge as the superior contenders, offering sharper physical accuracy, greater cross-scale adaptability, and enhanced numerical stability in the presence of strong nonequilibrium effects. Notably, although high-order models exhibit some degree of extensibility in capturing strong nonequilibrium effects, their applicability remains fundamentally constrained. In physical modeling, the key lies in dynamically adjusting the preserved kinetic moment relations during phase-space discretization and tailoring discretization strategies to specific research objectives-only then can we ensure a precise and robust kinetic description. To further unravel the multiscale nature of KHI, we construct a nonequilibrium phase diagram-a roadmap that visually charts the interplay between different nonequilibrium orders. This diagram not only deepens our understanding of the instability’s intricate dynamics but also serves as a valuable guide for selecting suitable kinetic models across a broad spectrum of nonequilibrium conditions in both scientific research and engineering applications.
@@ Appendix B Nonlinear constitutive relations
$\bm{\Delta}^{*(1)}_{2}=-\mu\left[\bm{\nabla}u+(\bm{\nabla}u)^{T}-\frac{2}{n+2}\mathbf{I}\bm{\nabla}\cdot\mathbf{u}\right],$ (48)